# A Fundamental Trade-off in Aligned Language Models and its Relation to Sampling Adaptors

### v2

@@ D.1 Prerequisites
${{\mathrm{H}}_{{\gamma}}}({p})=\begin{cases}\frac{1}{1-{\gamma}}\log{\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{p}({\boldsymbol{y}})^{\gamma}&{\gamma}\in(0,1)\\ -{\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{p}({\boldsymbol{y}})\log{p}({\boldsymbol{y}})&{\gamma}=1\end{cases}$
@@
A language model ${p}$ is eos-bounded if there exists $c$ such that ${p}({\textsc{eos}}\mid{\boldsymbol{y}})>c>0$ for all ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$. [^9] We note that for autoregressive language models, though tokens are sampled autoregressively from probability distributions over ${\overline{{{\Sigma}}}}$, the language model is a distribution over ${{{{\Sigma}}^{*}}}$.author=ryan,color=violet!40,size=,fancyline,caption=,]Rewrite this footnote. The point is that they are *not* equivalent. The is that ours is exact and theirs is an approximation. I called it “local renormalization”. In other words, ${p}({\boldsymbol{y}}\oplus{\textsc{eos}})\leq(1-c)^{|{\boldsymbol{y}}|}$ for all ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$.
@@ Appendix E Sampling Adaptors and String Probability
In this section, our goal is to show that an appropriate choice of globally normalized sampling adaptor can be used to control generated strings’ average probability under the prior ${p}({\boldsymbol{y}})$. Let us begin by decomposing a transform function ${\boldsymbol{\gamma}}$ into its constiutent components. Inspired by Meister et al. (2023a), we note that most ${\boldsymbol{\gamma}}$ can be formulated as the composition of truncation and scaling functions. The truncation function ${C}\colon{{{{\Sigma}}^{*}}}\rightarrow{\mathcal{P}}({\overline{{{\Sigma}}}})$ is a function used to find the set of tokens that meets specified criteria given the prior context, so that tokens deemed likely to lead to undesirable text can have their probability reassigned to other tokens, e.g., to only keep the top-${k}$ tokens. The scaling function ${\kappa}\colon{{\mathbbm{R}_{>0}}^{\mid{\overline{{{\Sigma}}}}\mid}}\rightarrow{{\mathbbm{R}_{>0}}^{\mid{\overline{{{\Sigma}}}}\mid}}$ is a simple scaling of the token probability, e.g., to the power of $\frac{1}{{\tau}}$ for some temperature parameter ${\tau}\in{\mathbbm{R}_{>0}}$. With these definitions we can express a transform function ${\boldsymbol{\gamma}}$ as
@@
That is, given a token distribution ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$, we apply the scaling function to scale token probabilities as needed and then remove tokens according to the truncation function to arrive at the output unnormalized distribution. For instance, we can express the transform function in nucleus sampling (Holtzman et al., 2020) with:
@@
Now, let us consider the probability of a string under the induced distribution ${{\widetilde{{p}}}_{+}}$ when a globally normalized sampling adaptor is used with an aligned model ${{p}_{{\textsc{+}}}}$. If we also apply Bayes’ rule, with some rearrangement we have:
${{\widetilde{{p}}}_{+}}({\boldsymbol{y}})$ $=\frac{{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({{y}_{t}})}{{{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}}$
$=\frac{\left({\mathbbm{1}}{\{{\textsc{eos}}\in{{C}_{|{\boldsymbol{y}}|}}\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\mathbbm{1}}{\{{{y}_{t}}\in{{C}_{t}}\}}\right)\left({\kappa}\big{(}{p}({\textsc{eos}}\mid{\boldsymbol{y}})\big{)}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\kappa}\big{(}{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}},{A}={\textsc{+}})\big{)}\right)}{{{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}}$ (41a)
@@
$\propto\left({\kappa}_{|{\boldsymbol{y}}|}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\kappa_{t}}\right)\left({\mathbbm{1}}{\{{\textsc{eos}}\in{{C}_{|{\boldsymbol{y}}|}}\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\mathbbm{1}}{\{{{y}_{t}}\in{{C}_{t}}\}}\right)\left({p}({\textsc{eos}}\mid{\boldsymbol{y}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}\right){p}({A}={\textsc{+}}\mid{\boldsymbol{y}})$ (41b)
@@
$=\underbrace{\left({\kappa}_{|{\boldsymbol{y}}|}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\kappa_{t}}\right)\left({\mathbbm{1}}{\{{\textsc{eos}}\in{{C}_{|{\boldsymbol{y}}|}}\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\mathbbm{1}}{\{{{y}_{t}}\in{{C}_{t}}\}}\right)}_{\text{Functions of the Prior}}{{p}_{{\textsc{+}}}}({\boldsymbol{y}})$ (41c)
where we use the shorthands ${\kappa_{t}}=\frac{{{\kappa}({p}(\cdot\mid{{\boldsymbol{y}}_{<t}}))({{y}_{t}})}}{{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}})}$, ${{C}_{t}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{C}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)}$ and ${{C}_{|{\boldsymbol{y}}|}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{C}\left({p}(\cdot\mid{\boldsymbol{y}})\right)}$. What Eq. 41 says is that by decomposing a transform function ${\boldsymbol{\gamma}}$ into its constitutent components, we see that the resulting probability of a string ${{\widetilde{{p}}}_{+}}({\boldsymbol{y}})$ is entirely determined by the choice of scaling and truncation functions, which take as input the probability of the string under the prior, i.e., ${p}({\boldsymbol{y}})$. As a result, an appropriate choice of globally normalized sampling adaptor can be used to effectively select strings based on their probability under the prior. [^11] So long as the transform functions can be expressed this way—such as those in § E.1. For example, using the transform function from top-${k}$ sampling will lead to the generation of corpora with higher average probability under the prior. [^12] For ${k}<|{\overline{{{\Sigma}}}}|$.
@@ E.1 Examples of Sampling Adaptors
i.e., the set of top-${k}$ most probable tokens.
@@
i.e., the set of items with log-probability closest to the token-level entropy that collectively have probability mass $\geq{\pi}$.
@@
that is, the set of tokens with probability greater than $\eta$, where $\eta=\min(\epsilon,\sqrt{\epsilon}\exp(-{\mathrm{H}}({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})))))$.
Appendix F The trade-off in DPO-aligned models

### v3

@@ D.1 Prerequisites
${{\mathrm{H}}_{{\gamma}}}({p})=\begin{cases}\frac{1}{1-{\gamma}}\log{\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{p}({\boldsymbol{y}})^{\gamma}&{\gamma}\in(0,1)\\ -{\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{p}({\boldsymbol{y}})\log{p}({\boldsymbol{y}})&{\gamma}=1\end{cases}$ (27)
@@
A language model ${p}$ is eos-bounded if there exists $c$ such that ${p}({\textsc{eos}}\mid{\boldsymbol{y}})>c>0$ for all ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$.111111 Strictly speaking, autoregressive language models are not always language models, i.e., valid probability distributions over ${{{{\Sigma}}^{*}}}$ where ${\sum_{{\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}}}{p}({\boldsymbol{y}})=1$. We highlight this difference in our exposition on locally vs. globally normalized sampling adaptors in § 4. In other words, ${p}({\boldsymbol{y}}\oplus{\textsc{eos}})\leq(1-c)^{|{\boldsymbol{y}}|}$ for all ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$.
@@ Appendix E Sampling Adaptors and String Probability
In this section, we provide, by means of a simple example, an intuition of how an appropriate choice of globally normalized sampling adaptor can be used to control generated strings’ average probability under the prior ${p}({\boldsymbol{y}})$. Inspired by Meister et al. (2023a), we note that most ${\boldsymbol{\gamma}}$ can be formulated as the composition of truncation and scaling functions. The truncation function ${C}\colon{{{{\Sigma}}^{*}}}\rightarrow{\mathcal{P}}({\overline{{{\Sigma}}}})$ is a function used to find the set of symbols that meets specified criteria given the prior context, so that symbols deemed likely to lead to undesirable text can have their probability reassigned to other symbols, e.g., to only keep the top-${k}$ symbols. The scaling function ${\kappa}\colon{{\mathbbm{R}_{>0}}^{\mid{\overline{{{\Sigma}}}}\mid}}\rightarrow{{\mathbbm{R}_{>0}}^{\mid{\overline{{{\Sigma}}}}\mid}}$ is a simple scaling of the symbol probability, e.g., to the power of $\frac{1}{{\tau}}$ for some temperature parameter ${\tau}\in{\mathbbm{R}_{>0}}$. With these definitions we can express a transform function ${\boldsymbol{\gamma}}$ as
@@
That is, given a symbol distribution ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$, we apply the scaling function to scale symbol probabilities as needed and then remove symbols according to the truncation function to arrive at the output unnormalized distribution. For instance, we can express the transform function in nucleus sampling (Holtzman et al., 2020) with:
@@
Let ${p}$ be a language model and ${{p}_{{\textsc{+}}}}$ its aligned counterpart. We consider the probability of a string under the induced distribution ${{\widetilde{{p}}}_{+}}$ when a globally normalized sampling adaptor is used with an aligned model ${{p}_{{\textsc{+}}}}$. Since our goal is simply to provide an intuition behind what might happen by analyzing the effects of the global sampling adaptor, we make the following simplifying assumptions. First, we assume that the alignment does not modify the truncation function ${C}$: ${C}\left({p}\left(\cdot\mid{\boldsymbol{y}}\right)\right)={C}\left({{p}_{{\textsc{+}}}}\left(\cdot\mid{\boldsymbol{y}}\right)\right)$ for all ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$. Second, we assume that ${\kappa}\left(\cdot\right)={\mathbb{I}}\left(\cdot\right)$. With this, we can derive
${{\widetilde{{p}}}_{+}}({\boldsymbol{y}})$ $\propto{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{{\boldsymbol{y}}_{<t}})\big{)}({{y}_{t}})$
@@
$={\kappa}({{p}_{{\textsc{+}}}}({\textsc{eos}}\mid{\boldsymbol{y}})){\mathbbm{1}}{\{{\textsc{eos}}\in{C}\left({{p}_{{\textsc{+}}}}\left(\cdot\mid{\boldsymbol{y}}\right)\right)\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\kappa}({{p}_{{\textsc{+}}}}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}})){\mathbbm{1}}{\{{{y}_{t}}\in{C}\left({{p}_{{\textsc{+}}}}\left(\cdot\mid{{\boldsymbol{y}}_{<t}}\right)\right)\}}$ (42a)
$={{p}_{{\textsc{+}}}}({\textsc{eos}}\mid{\boldsymbol{y}}){\mathbbm{1}}{\{{\textsc{eos}}\in{C}\left({{p}_{{\textsc{+}}}}\left(\cdot\mid{\boldsymbol{y}}\right)\right)\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{{p}_{{\textsc{+}}}}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}){\mathbbm{1}}{\{{{y}_{t}}\in{C}\left({{p}_{{\textsc{+}}}}\left(\cdot\mid{{\boldsymbol{y}}_{<t}}\right)\right)\}}$ (42b)
$={{p}_{{\textsc{+}}}}({\textsc{eos}}\mid{\boldsymbol{y}}){\mathbbm{1}}{\{{\textsc{eos}}\in{C}\left({{p}_{{\textsc{+}}}}\left(\cdot\mid{\boldsymbol{y}}\right)\right)\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{{p}_{{\textsc{+}}}}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}){\mathbbm{1}}{\{{{y}_{t}}\in{C}\left({{p}_{{\textsc{+}}}}\left(\cdot\mid{{\boldsymbol{y}}_{<t}}\right)\right)\}}$ (42c)
$={{p}_{{\textsc{+}}}}({\textsc{eos}}\mid{\boldsymbol{y}}){\mathbbm{1}}{\{{\textsc{eos}}\in{C}\left({p}\left(\cdot\mid{\boldsymbol{y}}\right)\right)\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{{p}_{{\textsc{+}}}}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}){\mathbbm{1}}{\{{{y}_{t}}\in{C}\left({p}\left(\cdot\mid{{\boldsymbol{y}}_{<t}}\right)\right)\}}$ (42d)
$={\mathbbm{1}}{\{{\textsc{eos}}\in{C}\left({p}\left(\cdot\mid{\boldsymbol{y}}\right)\right)\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\mathbbm{1}}{\{{{y}_{t}}\in{C}\left({p}\left(\cdot\mid{{\boldsymbol{y}}_{<t}}\right)\right)\}}{{p}_{{\textsc{+}}}}({\textsc{eos}}\mid{\boldsymbol{y}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}})$ (42e)
$={\mathbbm{1}}{\{{\textsc{eos}}\in{C}\left({p}\left(\cdot\mid{\boldsymbol{y}}\right)\right)\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\mathbbm{1}}{\{{{y}_{t}}\in{C}\left({p}\left(\cdot\mid{{\boldsymbol{y}}_{<t}}\right)\right)\}}{{p}_{{\textsc{+}}}}({\boldsymbol{y}})$ (42f)
@@
$\propto\underbrace{{\mathbbm{1}}{\{{\textsc{eos}}\in{C}\left({p}\left(\cdot\mid{\boldsymbol{y}}\right)\right)\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\mathbbm{1}}{\{{{y}_{t}}\in{C}\left({p}\left(\cdot\mid{{\boldsymbol{y}}_{<t}}\right)\right)\}}}_{F_{\boldsymbol{\gamma}}\left[{{p}({\boldsymbol{y}})}\right]}{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}\left({\boldsymbol{y}}\right)\right)$ (42g)
$=F_{\boldsymbol{\gamma}}\left[{{p}({\boldsymbol{y}})}\right]{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}\left({\boldsymbol{y}}\right)\right)$ (42h)
We see that, in this simplified setting, the resulting probability of a string under the adapted and aligned model ${{\widetilde{{p}}}_{+}}$ equals the prior probability scaled by the reward function and by the truncation factors. Since different sampling adaptors affect this relationship differently, this simplified example suggests that an appropriate choice of globally normalized sampling adaptor can be used to effectively select strings based on their probability under the prior. For example, using the transform function from top-${k}$ sampling will lead to the generation of corpora with higher average probability under the prior, [^13] For ${k}<|{\overline{{{\Sigma}}}}|$. and thus, by the anti-correlation derived in our paper, with lower average reward.
@@ E.1 Examples of Sampling Adaptors
i.e., the set of top-${k}$ most probable symbols.
@@
i.e., the set of items with log-probability closest to the symbol-level entropy that collectively have probability mass $\geq{\pi}$.
@@
that is, the set of symbols with probability greater than $\eta$, where $\eta=\min(\epsilon,\sqrt{\epsilon}\exp(-{\mathrm{H}}({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})))))$.
Appendix F The Probability–Quality Trade-off in DPO-aligned Language Models